, Break-Even Sales Revenue The c Sales Total Variable cost $88,000 68,640 $19,360 9,680 $9,680 Contribution margin Total Fixed cost Operating income Required: 1. Calculate the contribution margin ratio 2. Caiculate the variable cost ratio 3. Calculate the break-even sales revenue for Ashton. the total sales revenue? 4. How could Ashton increase projected operating income without in

media%2F2c6%2F2c69ee8c-d1cb-4944-bc5c-eb

media%2F2b7%2F2b7b52ab-815b-43cb-8968-96
Answer #1

1a) Contribution margin ratio = 19360/88000 = 22%

1b) Variable expense ratio = 100-22 = 78%

1c) Break even sales = 9680/.22 = $44000

1d) By reducing fixed cost

Note : Please post each question individually as HOMEWORKLIB RULES guidelines